### v4

# Quantum-Mechanical Suppression of Accretion by Primordial Black Holes

Abstract
The Schwarzschild radii of primordial black holes (PBHs) in the mass range of $6\times 10^{14}~{}{\rm g}$ to $4\times 10^{19}~{}{\rm g}$ match the sizes of nuclei to atoms. I discuss the resulting quantum-mechanical suppression in the accretion of matter by PBHs in dense astrophysical environments, such as planets or stars.
1 Introduction
Current cosmological constraints allow for the possibility that dark matter is made of primordial black holes (PBHs) in the asteroid mass range of $\sim 10^{18}$-$10^{22}~{}{\rm g}$ (Carr & Hawking, 1974; Carr & Kuhnel, 2021; Green, 2024; Carr & Green, 2024).
The accretion of atomic gas or a plasma onto PBHs was described so far (De Luca et al., 2022; De Luca & Bellomo, 2023; Fowler & Anantua, 2023; Curd et al., 2024) based on the hydrodynamics of a continuous fluid. Here, I point out that this formalism is invalid because of quantum-mechanical effects which dominate for asteroid-mass PBHs.
A classical black hole is the ultimate prison. However, it is difficult to fit a plump prisoner into a prison that is much smaller than the prisoner’s body size. In the case of a black hole, the prison walls are represented by the event horizon, which in the non-spinning case is a sphere with the Schwarzschild radius,
$r_{\rm Sch}=\left({2GM\over c^{2}}\right)=5.3\times 10^{-9}~{}{\rm cm}\left({M\over 3.6\times 10^{19}~{}{\rm g}}\right)~{},$ (1)
where $G$ is Newton’s constant, $c$ is the speed of light and $M$ is the black hole mass (Schwarzschild, 1916). The Schwarzschild radius equals the Bohr radius of the hydrogen atom (Bohr, 1913),
$r_{\rm B}={\hbar^{2}\over m_{e}e^{2}}=5.3\times 10^{-9}~{}{\rm cm}~{},$ (2)
for a black hole mass of $M=3.6\times 10^{19}~{}{\rm g}$. PBHs of this mass could have been produced in the early Universe. The Schwarzschild radius equals the radius of the proton, $r_{p}=8.4\times 10^{-14}~{}{\rm cm}$ (Antognini et al., 2013) for a black hole mass of $M=5.7\times 10^{14}~{}{\rm g}$. This happens to be a few times the mass of a PBH that evaporates by Hawking radiation (Hawking, 1974) on a timescale comparable to the age of the Universe. Much smaller PBHs would have disappeared by now.
2 Quantum Suppression of Astrophysical Accretion
Black holes grow in mass by accreting matter from their astrophysical environment (Yuan & Narayan, 2014; Blaes, 2014). For black holes more massive than the Sun, the event horizon is large enough to be considered as a mouth that absorbs many atoms at once and so the accreted matter can be approximated as a continuous fluid. But in the regime of asteroid-mass PBHs, this assumption is no longer valid. These black holes feast on a single atom, a single proton or a single electron at a time, because their event horizon is smaller than the spatial extent of the quantum-mechanical wave function of these particles.
Consider a situation where a hydrogen atom is attracted gravitationally towards a PBH. In the naive perception of the atom as a point particle, it could reach the black hole center from a distance $r$ over a free-fall time, $\sim(r^{3}/GM)^{1/2}$. However, if the horizon is smaller than the size of the atom, then this ‘plump prisoner’ will not be captured and most of the atom will stay predominantly outside the event horizon. Even if the PBH is massive enough to absorb the more compact proton, the extended wave function of the electron could remain outside the PBH. In that case, the PBH will acquire a positive electric charge. The electric force that binds the electron to the proton will be enhanced by the PBH gravity.
At distances much larger than the horizon, both electric and gravitational forces decline inversely with distance squared. Hence, the effective Bohr radius (marked hereafter by a tilde) of the hydrogen$+$PBH atom shrinks to a value,
${\tilde{r}}_{\rm B}={\hbar^{2}\over m_{e}(e^{2}+GMm_{e})}=5.3\times 10^{-9}~{}{\rm cm}\times\left[1+\left({M\over 3.8\times 10^{15}~{}{\rm g}}\right)\right]^{-1}~{}.$ (3)
Correspondingly, the effective Rydberg energy levels will be modified to $E_{n}=-1{\rm{\tilde{R}y}}/n^{2}$, with $n$ an integer. The modified binding energy of the ground state would be,
$-E_{1}=1{\rm{\tilde{R}y}}={\hbar^{2}\over 2m_{e}{\tilde{r}}_{\rm B}^{2}}=13.6~{}{\rm eV}\times\left[1+\left({M\over 3.8\times 10^{15}~{}{\rm g}}\right)\right]^{2}~{}.$ (4)
The gravitational and electric forces are of equal magnitude for a PBH mass of $M=3.8\times 10^{15}~{}{\rm g}$, and gravity wins at larger masses. The atomic radius in equation (3) shrinks below the Schwarzschild radius in equation (1) for a PBH mass $M=3.7\times 10^{17}~{}{\rm g}$. At this mass, the gravitational binding energy in equation (4) approaches a quarter of the electron rest-mass energy. For higher PBH masses, the Compton wavelength of the electron, $(\hbar/m_{e}c)=4\times 10^{-11}~{}{\rm cm}$, is smaller than the Schwarzschild radius, and the ground state of the bound electron can be inside the horizon. The criterion that accretion of an electron is suppressed is identical, up to a numerical factor of order unity, to the condition that electron-positron pairs are emitted through Hawking evaporation of the black hole.
For lower mass PBHs, accretion could be significantly suppressed in common astrophysical environments, where gas particles have non-relativistic thermal speeds and the mass density is much lower than the nuclear density of neutron stars (Giffin et al., 2022). In the rarefied environments of the interstellar or intergalactic media, a low-mass PBH would accrete one particle at a time with a significant quantum-mechanical suppression in accretion rate relative to hydrodynamic expectations.
3 Accretion Rate
To obtain reliable assessments of the accretion rate by PBHs with masses $M\lesssim 4\times 10^{19}~{}{\rm g}$, one must calculate the wave function of electrons and protons using the Dirac equation in the background PBH metric. For each quantum state and energy level, the overlap of the electron wave function with the volume interior to the event horizon sets a finite half-life for the electron to stay in a bound state outside the horizon. Afterwards, the electron will join the proton inside the horizon and neutralize the PBH charge.
In the quantum world, there is a finite probability per unit time for a plump atom to be captured by a small event horizon. The quantum transition to the final state of capture resembles tunneling through a barrier.
3.1 Rarefied Cold Gas with Bound Electrons
The accretion of particles in bound states around a PBH or in tightly packed systems is different from absorption of free particles by a PBH (Unruh, 1976). A bound electron can be captured by an atomic nucleus even without a central black hole. This well-known process, called electron capture, involves the absorption an electron from an inner atomic shell by a proton-rich nucleus of a neutral atom (Alvarez, 1937). The rate for this capture is related to the overlap of the electron wave function with the volume of the atomic nucleus.
Similarly, the accretion of protons and electrons by a PBH in the asteroid mass range, is dictated by the overlap of their quantum-mechanical wave function with the volume interior to the PBH event horizon.
Ignoring Hawking radiation, an estimate for the half-life of a bound electron around a black hole with $r_{\rm Sch}<{\tilde{r}}_{\rm B}$ (or equivalently $M<3.7\times 10^{17}~{}{\rm g}$) which already captured a proton, is
$\tau_{1/2}\sim{({\tilde{r}}_{\rm B}/{\tilde{v}}_{\rm B})\over(r_{\rm Sch}/{\tilde{r}}_{\rm B})^{3}}=2.2\times 10^{-11}~{}{\rm s}\left({M\over 3.8\times 10^{15}~{}{\rm g}}\right)^{-3}\left[1+\left({M\over 3.8\times 10^{15}~{}{\rm g}}\right)\right]^{-5}~{},$ (5)
where the effective electron speed is,
${\tilde{v}}_{\rm B}\sim\left[{(e^{2}+GMm_{e})\over\hbar c}\right]c=2.2\times 10^{8}~{}{\rm cm~{}s^{-1}}\left[1+\left({M\over 3.8\times 10^{15}~{}{\rm g}}\right)\right]~{}.$ (6)
The accretion rate corresponding to the absorption of an electron-proton pair per $\tau_{1/2}$ is,
${\dot{M}}\equiv{m_{p}\over\tau_{1/2}}=7.5\times 10^{-14}~{}{\rm g~{}s^{-1}}~{}\left({M\over 3.8\times 10^{15}~{}{\rm g}}\right)^{3}\left[1+\left({M\over 3.8\times 10^{15}~{}{\rm g}}\right)\right]^{5}~{}.$ (7)
The accretion rate is too low to add significant mass to PBHs with $M\lesssim 1.4\times 10^{17}~{}{\rm g}$ over the entire age of the Universe.
At high enough infall rate, a bound state could involve multiple electrons and protons simultaneously. When the inflow of fresh protons into the PBH exceeds $\tau_{1/2}^{-1}$, the PBH could be charged positively up to a maximum charge, $Q_{\rm max}\sim(k_{\rm B}T/m_{p})(GMm_{p}/e^{2})e=(T/1.1\times 10^{13}~{}{\rm K})(M/2.1\times 10^{12}~{}{\rm g})e$, at which the electric potential near the horizon equals the thermal energy for external protons at the background temperature $T$ (Medvedev & Loeb, 2024). At this maximum charge (which will be screened over the Debye length in the surrounding plasma), the accretion rate can be enhanced relative to equation (7) since the effective Bohr radius for the ground state of the innermost electron at $Q_{\rm max}$ shrinks below the Schwarzschild radius for PBH masses $M\gtrsim 8.6\times 10^{15}~{}{\rm g}$. However, at the high plasma temperatures expected near the PBH horizon, the electrons will not stay in bound states and the accretion rate will be dictated by the dynamics of free particles.
3.2 Dense Hot Plasma with Free Electrons
For a thermal plasma of free (unbound) particles, the accretion rate can exceed the value expected from equation (7) if the interparticle distance $d=(4\pi n_{e}/3)^{-1/3}$ is shorter than ${\tilde{r}}_{\rm B}$, where $n_{e}$ is the electron number density in the plasma. In this regime, the fastest accretion will involve one electron-proton pair per $\tau_{1/2}$, where the length scale ${\tilde{r}}_{\rm B}$ in equation (5) is replaced by the interparticle distance $d$ and the velocity scale ${\tilde{v}}_{\rm B}$ is replaced by the thermal speed of the particles $v_{\infty}$ far from the PBH (assuming a stationary PBH relative to the plasma rest frame). This substitution gives a net accretion rate of,
${\dot{M}}={4\over 3}\left({r_{\rm Sch}\over d}\right)\rho_{\infty}v_{\infty}\left(\pi r_{\rm Sch}^{2}\right)~{},$ (8)
where $\rho_{\infty}=m_{p}n_{e}$ is the plasma mass density far from the PBH.
It is instructive to compare this suppressed accretion rate to the Bondi accretion rate (Bondi, 1952), derived in the fluid approximation,
${\dot{M}}_{\rm Bondi}\approx\left({c\over v_{\infty}}\right)^{2}\rho_{\infty}v_{\infty}\left(\pi r_{\rm Sch}^{2}\right)~{}.$ (9)
The resulting suppression factor for $d>r_{\rm Sch}$ is,
$F_{\rm supp}=\left({{\dot{M}}\over{\dot{M}}_{\rm Bondi}}\right)\approx\left({v_{\infty}\over c}\right)^{2}\left({4\over 3}{r_{\rm Sch}\over d}\right)~{}.$ (10)
For $r_{\rm Sch}\ll d$ and $v_{\infty}\ll c$, we find that $F_{\rm supp}\ll 1$.
This result applies to high matter densities at which atoms are densely packed and $d<\tilde{r}_{\rm B}$. Such densities are relevant when PBHs pass through the interiors of planets or stars (Hawking, 1971; Luo et al., 2012; Caplan et al., 2024), with the most dense medium realized for PBHs trapped in the interior of a neutron star (Pani & Loeb, 2014; Giffin et al., 2022). For context, the Bondi accretion rate of a trapped PBH at the mean density of a white dwarf is substantial for $v_{\infty}<5\times 10^{3}~{}{\rm km~{}s^{-1}}$,
${\dot{M}}_{\rm Bondi}\approx 1.5\times 10^{5}~{}{\rm g~{}s^{-1}}\left({\rho_{\infty}\over 10^{6}~{}{\rm g~{}cm^{-3}}}\right)\left({M\over 3.6\times 10^{19}~{}{\rm g}}\right)^{2}\left({v_{\infty}\over 5\times 10^{3}~{}{\rm km~{}s^{-1}}}\right)^{-3}.$ (11)
Expressing this equation as ${\dot{M}}=AM^{2}$ with $A$ being a constant, yields a solution for the final PBH mass $M_{f}$ in terms of its initial mass $M_{i}$ as, $M_{f}=1/(1/M_{i}-At)$. This solution implies that over a timescale shorter than 10 Myr, the mass of a trapped PBH with $M_{i}=3.6\times 10^{19}~{}{\rm g}$, would diverge through Bondi accretion well above its initial value. Consequently, a trapped PBH could have had dramatic observable consequences on the energy budget or the ultimate fate of a white dwarf. However, in reality the PBH growth would be suppressed by the quantum-mechnical effects discussed above as well as by the Eddington limit on the accretion rate, which is $\sim 3\times 10^{5}~{}{\rm g~{}s^{-1}}(\epsilon/0.01)^{-1}(M/3.6\times 10^{19}~{}{\rm g})$, where $\epsilon$ is the radiative efficiency.
Other dense environments involve the thermal electron-positron plasma of the early universe (Curd et al., 2024), where the thermal speed of the particles is $\sim c$. The de Broglie wavelength of relativistic electrons and positrons in this thermal plasma of free particles is of order $d$, smaller than their Compton wavelength. This is properly incorporated in the suppression factor of equation (10) which is effective for $d>r_{\rm Sch}$. For a wavefunction with a de Broglie wavelength larger than the Schwarzschild radius, the relevant momentum is calculated far from the horizon, where most of the probability distribution of finding the particle resides.
4 Concluding Remarks
Including Hawking radiation (Hawking, 1974) would further lower the accretion rate since the Hawking temperature, $T_{\rm H}=(\hbar c^{3}/8\pi k_{B}GM)=2.7\times 10^{6}~{}{\rm eV}~{}(M/3.8\times 10^{15}~{}{\rm g})^{-1}$, exceeds the binding energy $1{\rm{\tilde{R}y}}$ in equation (4), and the Hawking luminosity, $L_{\rm H}=(\hbar c^{6}/15360\pi G^{2}M^{2})=1.5\times 10^{26}~{}{\rm eV~{}s^{-1}}~{}(M/3.8\times 10^{15}~{}{\rm g})^{-2}$, exceeds $1{\rm{\tilde{R}y}}/\tau_{1/2}$. The outward flux of high energy photons and electron-positron pairs could suppress accretion altogether.
The accretion of matter into asteroid-mass PBHs is only significant in the dense interiors of planets or stars, but can be ignored at the much lower cosmic-mean-density of matter and radiation in the Universe (Rice & Zhang, 2017).
If we ever witness an asteroid-mass black hole in the solar system, it could serve as a testbed for quantum-gravitational physics on a subatomic scale.
Acknowledgements
I am grateful for insightful comments from an anonymous referee. This work was supported in part by Harvard’s Black Hole Initiative, which is funded by grants from JFT and GBMF.